|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0227smov.html
Дата изменения: Tue Mar 11 23:59:06 1997 Дата индексирования: Thu Feb 27 23:54:13 2014 Кодировка: Поисковые слова: http www.badastronomy.com phpbb index.php |
STATUS REPORTS
1. Observatory Status
MOSES status reports:
-PCS - nominal
-OTA - nominal
-EPS - nominal
-DMS - nominal
- SSR testing, including SAA behavior,
continues in data gathering mode with recurrence of
now typical re-tries.
-Thermal - nominal.
2. WFPC2 Status - nominal
- CCD temperature now at -88C
- Today's UV monitor's at this temperature will be
used to decide whether or not to remain
at the operational -88C. A decision is expected
at around 10 pm tonight.
3. NICMOS Status - nominal
The cold well temp is at 56K and levelling off.
VCS temperature is within the expected range.
NICMOS science is in the process of analyzing data from
recent internal tests. (Note: Since this meeting, NICMOS
went into suspend with a Detector 1 VD voltage
out of limits low. Reason unknown, more details in
tomorrow's meeting notes.)
4. STIS Status - CCD temperature is nominal -83 degC. MAMA low
voltages were turned back on yesterday, and since that time
another MCE reset, this time on MAMA1, has occurred, again
in the SAA. The plan is to reset (or unreset) the affected
parameters and keep running. (Note: Since this meeting,
a third MCE reset (MAMA1) has occurred. More details
in tomorrows notes and from the planned Project meeting
this afternoon.)
5. FGS status - no report
6. FOC/COSTAR status - nominal.
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS058 - Currently executing.
- SMS062 - signed off and ready for uplink.
- SMS065 - finals are out for review
- SMS069 - built and initials out for review
- SMS072 - built and initials out for review
- SMS076 - built
- SMS079 - being worked
2. Replans - none
OPERATIONS STATUS -
1. OPUS
- OMS data is now being processed by through the pipeline.
- A problem with the FITS I/O kernel for CALSTIS has been
identified and a fix is expected for next week. In the
meantime, any STIS science data that fails calibration
will be FTP'd directly to the team without going through
CALSTIS.
2. Archive - A problem with an OMS STSDAS task has been identified
and is being worked.